OpenStudy (anonymous):

Perform the indicated operation. Write simplest form. 1/6 x [5 3/4 - 1/5]

OpenStudy (anonymous):

basic fraction// learn fractions dear!

OpenStudy (anonymous):

its 5 X 3 or 53?

OpenStudy (anonymous):

it is 1 ( 23 1 ) 1 115 - 4 1 111 ---- x (---- - -----) = ---- x ---------- = ----- x ------ 6 ( 4 5 ) 6 20 6 20 now do the rest n get yr ans

OpenStudy (anonymous):

\[\frac{1}{6} \times (\frac{23}{4} -\frac{1}{5})\rightarrow\frac{1}{6}\times\frac{111}{20}=\frac{111}{120}\]